OpenStudy (anonymous):

Suppose a fuel mixture is 7% ethanol and 93% gasoline. How much ethanol (in gallons) must you add to one gallon of the fuel so that the new fuel mixture is 10% ethanol?

OpenStudy (anonymous):

\[.07+x=.10(1+x)\]

OpenStudy (anonymous):

you have one gallon to begin with, which is 7% ethanol, so it has .07 of a gallon of ethanol. if you add x amount of ethanol you will have .07 + x amount of ethanol and \[1+x\] amount of mixture. you want this to be 10% ethanol, so you want \[.07+x=.10(1+x)\]

OpenStudy (anonymous):

easier to solve if you write \[7+100x=10(1+x)\] i.e. multiply everything by 100 then you get \[7+100x=10+10x\] \[90x=3\] \[x=\frac{3}{90}=\frac{1}{30}\]